OpenStudy (anonymous):

Solve for y and use a graphing utility to graph each of the resulting equations in the same viewing window.

OpenStudy (anonymous):

\[(x+2)^2 + (y-3)^2 = 25\]

OpenStudy (anonymous):

Adjust the viewing window so that a circle really does appear circular.

OpenStudy (anonymous):

5000 x the sqrt of 55 = PONIES :P

OpenStudy (anonymous):

(y-3)^2=25-(x+2)^2, \[y-3=\sqrt{(x+2)^{2}+25}\], \[y=3+\sqrt{(x+2)^{3}+25}\]

OpenStudy (anonymous):

come at me bro

OpenStudy (anonymous):

What? NO!

OpenStudy (anonymous):

Can I ask why you put it under a square root?

OpenStudy (anonymous):

Cos he's a boss like that...U MAD BRO?

OpenStudy (anonymous):

to get rid of the square in (y-3)^2

OpenStudy (anonymous):

Oh, gotcha. Thanks! Appreciate it.

OpenStudy (anonymous):

we coooooool brah